OpenStudy (anonymous):

The product of twenty less than eight times a number and thirty less than two times a number equals zero, What are the numbers?

OpenStudy (campbell_st):

well they are binomials... \[(8x - 20)\times (2y - 30) = 0\] it just means that either 8x - 20 = 0 or 2y - 30 = 0 this will result in one binomial being zero... just solve for x and y

OpenStudy (anonymous):

how do i solve for x and y ?

OpenStudy (campbell_st):

well 8x = 20 divide by x 2y = 30 divide by 2

OpenStudy (anonymous):

so thw answer would be 15 ?

OpenStudy (anonymous):

& 2.5

OpenStudy (campbell_st):

yes... i think that works...

OpenStudy (anonymous):

yeah thank you (: